How much water must be evaporated from 32 ounces of a 2% salt solution to make a 8% salt solution?
Answer by robertb(5830)

0.02*32 - 0.00*x = 0.08(32 - x)
==> 0.64 = 2.56 - 0.08x
==> 0.08x = 1.92
==> x = 24 ounces of water must be evaporated.